### 05 Mar 2026 v9

# Efficient frontiers for portfolios under SSD and law-invariant risk measures with hyperbolic return distributions

Abstract
The classical Markowitz mean–variance framework treats variance as the measure of risk and determines efficient frontier portfolios in closed form using standard optimization methods. For more general mean–risk settings, however, obtaining closed-form optimal portfolios is typically difficult. In this note, we derive explicit expressions for frontier portfolios corresponding to arbitrary law-invariant convex risk measures, assuming that the return vector follows a normal mean–variance mixture distribution. Our approach begins by developing stochastic dominance conditions within the family of normal mean–variance mixture models, which we then utilize to derive closed-form representations of frontier portfolios. The central result establishes that if asset returns are described by a normal mean–variance mixture model, then the associated mean–risk efficient frontier can be obtained by solving a Markowitz mean–variance problem with a suitably modified return vector. Moreover, we show that for any normal mean–variance mixture specification of the form $\eta=a+bZ+c\sqrt{Z}N$, the following relationship holds under every law-invariant coherent risk measure
$\rho(\eta)=a+bEZ+c\rho(\sqrt{Z}N).$
We then use this result to obtain a CAPM-style equilibrium model within our framework.
Keywords: Frontier portfolios; Mean-variance mixtures; Risk measures; Stochastic dominance; Mean-risk criteria
JEL Classification: G11
1 Motivation
In risk management and portfolio optimization, it is crucial to describe the underlying asset returns with appropriate probability distributions. The traditional Markowitz portfolio optimization framework relies on the assumption that asset returns follow a normal distribution. However, this normality assumption has been extensively challenged in the literature. A large body of empirical work on asset returns shows that models based on normality are often inconsistent with observed data. As the foundational study [6] illustrates, empirical return densities exhibit higher concentration around the mean and in the tails, and lower mass in the intermediate regions, than the normal distribution.
Another important property of empirical asset returns is their asymmetry, typically captured by skewness. Skewness in asset returns was identified early in the evolution of modern finance theory, and since then substantial work has focused on developing more refined models for return dynamics, motivated by the belief that improved models would enable better financial decisions. Consequently, it is clear that robust financial decision-making must be grounded in a well-specified multivariate probability distribution for asset returns.
Generalized hyperbolic distributions constitute a versatile family that can replicate most empirically observed characteristics of asset returns. As demonstrated in [6], hyperbolic distributions achieve an almost exact match to observed return density functions. Hyperbolic distributions are, in turn, a special case within the more general class of Normal mean-variance mixture (NMVM) models.
A $d$-dimensional random vector $X$ is said to follow a normal mean–variance mixture distribution with mixing variable $Z$ if, for each fixed $Z=z\in\mathbb{R}_{+}=(0,+\infty)$, the conditional distribution of $X$ is $X\mid_{Z=z}\sim W_{d}(\mu+z\gamma,z\Sigma)$. Here, $\mu,\gamma\in\mathbb{R}^{d}$ are fixed vectors, $\Sigma$ is a fixed positive-definite $d\times d$ matrix, and $W_{d}$ denotes a $d$-dimensional normal random vector with mean $\mu+z\gamma$ and covariance matrix $z\Sigma$. It is well known that such a random vector $X$ can be represented as
$X\overset{d}{=}\mu+\gamma Z+\sqrt{Z}\,AN_{d},$ (1)
where $A=\Sigma^{1/2}$ and $N_{d}$ is a $d$-dimensional normal random vector with mean equal to the zero vector in $\mathbb{R}^{d}$ and covariance matrix given by the identity matrix $I_{d}$ in $\mathbb{R}^{d}\times\mathbb{R}^{d}$. The mixing variable $Z$ in (1) is assumed independent of $N_{d}$.
Within the framework of (1), several specifications have become particularly popular in financial modeling. When $Z$ is assumed to follow a Generalized Inverse Gaussian (GIG) distribution, the resulting distribution of $X$ is a multi-dimensional generalized hyperbolic (mGH) distribution. This family of distributions was originally introduced in [2], and their extensive applications in finance have been explored in a number of subsequent works, including [1], [9], [6], [22], [21], [8], and [7]. For instance, [22] shows that the variance-gamma model and its elliptical multivariate extension provide an excellent description of asset returns over a unit time horizon. Similarly, [6] demonstrates that hyperbolic distributions can deliver an almost perfect fit to empirical return density functions. In [1], the authors find that the Generalized Hyperbolic Skew Student’s $t$ distribution matches empirical financial data extremely well.
In the multivariate context, [24] calibrates the multivariate generalized hyperbolic (mGH) model to both multivariate stock returns and multivariate exchange rate returns and shows, via likelihood-ratio tests, that the Gaussian specification is consistently rejected in favor of the more general mGH class. Furthermore, [1] employs multivariate NIG (Normal Inverse Gaussian) distributions in a risk management setting and demonstrates that they fit the empirical distribution of hedge fund returns substantially better than the Normal distribution. Generalized hyperbolic distributions have also been used in the valuation of interest rate derivatives, where they have been shown to yield highly accurate prices for caplets and other interest rate products; see, for example, [8] and [7].
In this paper, we examine a financial market consisting of $d$ risky assets and assume that their return vector $X$ satisfies (1). The set of admissible portfolios is $\mathbb{R}^{d}$, and for any portfolio $\omega\in\mathbb{R}^{d}$ the associated portfolio return is $\omega^{T}X$, where $T$ denotes transposition. The central objective of this work is to analyze the optimization problem
[…]
for any prescribed target level of expected return $r\in\mathbb{R}$. In (2), $\rho$ represents a risk measure and $\mathbf{1}$ denotes the $d$-dimensional vector whose components are all equal to one. In this note, we derive explicit closed-form solutions to (2) for a class of risk measures $\rho$ satisfying certain assumptions that will be specified in Section 3.
One of these assumptions is that $\rho$ is law-invariant, meaning that $\rho(\eta_{1})=\rho(\eta_{2})$ whenever $\eta_{1}\overset{d}{=}\eta_{2}$. Using the relation
$\omega^{T}X\overset{d}{=}\omega^{T}\mu+\omega^{T}\gamma Z+\sqrt{Z}\omega^{T}\Sigma\omega N(0,1),$
where $N(0,1)$ is a standard normal random variable independent of $Z$, it follows that
$\rho(-\omega^{T}X)=\rho(-\omega^{T}\mu-\omega^{T}\gamma Z+\sqrt{Z}\omega^{T}\Sigma\omega N(0,1)).$
We further impose that $\rho$ is compatible with second-order stochastic dominance. Consequently, as will become clear, the analysis of problem (2) can be reduced to studying stochastic dominance relations among one-dimensional NMVM models.
We work on an atomless probability space $(\Omega,\mathcal{F},P)$, on which all random variables in this paper are defined. We write $L^{0}=L^{0}(\Omega,\mathcal{F},P)$ for the set of all real-valued random variables that are almost surely finite, and for $1\leq k<\infty$ we denote by $L^{k}=L^{k}(\Omega,\mathcal{F},P)$ the space of random variables with finite $k$-th moment. The space $L^{\infty}(\Omega,\mathcal{F},P)$ consists of all essentially bounded random variables. For vectors $\mathbf{x}=(x_{1},\dots,x_{d})^{T}$ and $\mathbf{y}=(y_{1},\dots,y_{d})^{T}$, we denote their scalar product by $\mathbf{x}\cdot\mathbf{y}=\mathbf{x}^{T}\mathbf{y}=\sum_{i=1}^{d}x_{i}y_{i},$ and the Euclidean norm of $\mathbf{x}$ by $|\mathbf{x}|=(x_{1}^{2}+\cdots+x_{d}^{2})^{1/2}.$ Following [10], for each positive integer $d$ we set $L_{d}^{k}=(L^{k})^{d}$, the space of $d$-dimensional random vectors whose components all lie in $L^{k}$. Unless stated otherwise, we assume throughout that the mixing variable $Z$ belongs to $L^{1}$. We additionally suppose that $Z$ has a probability density function $f_{Z}(z)$. This latter assumption is not required for the theoretical results of the paper, but it is imposed to streamline the calculations. We denote by $\phi(\cdot)$ the probability density function of a standard normal random variable and by $\varphi(\cdot)$ its cumulative distribution function.
Our original motivation for investigating problem (2) was to analyze its solutions under the widely used and popular risk measures $VaR$ and $CVaR$. The value-at-risk (VaR) is defined for any random variable in $L^{0}$. For $\eta\in L^{0}$, we denote by $q_{\alpha}^{-1}(\eta)=\inf\{a\in\mathbb{R}:F_{\eta}(a)\geq\alpha\}$ the left-continuous inverse of the cumulative distribution function $F_{\eta}(\cdot)$ of $\eta$. For each $\alpha\in(0,1)$, the quantity $VaR_{\alpha}$ is then given by $VaR_{\alpha}(\eta)=q_{\alpha}^{-1}(\eta)$. The conditional value-at-risk (CVaR) is defined on $L^{1}$ by $CVaR_{\alpha}(\eta)=(1/(1-\alpha))\int_{\alpha}^{1}q_{\beta}(\eta)d\beta$ for any $\eta\in L^{1}$.
In this note, we are able to obtain a closed-form solution to (2) when $\rho=CVaR_{\alpha}$, whereas deriving an analytical solution to (2) when $\rho=VaR_{\alpha}$ appears to be difficult. The main reason is that the risk measure $CVaR_{\alpha}$ is compatible with second order stochastic dominance. Other related property of CVaR is its continuity on $L^{k}$ for every $k\geq 1$; see Theorem 4.1 in [17] and page 11 of [10] and the references cited therein. This property of CVaR is not explicitly utilized in the computations presented in our paper. Nevertheless, the proof of Proposition 3.12 could alternatively be derived by using CVaR in place of the method adopted in the current proof of the proposition. Moreover, although we do not use it here, it is also known that CVaR is Lipschitz continuous with respect to $L^{k}$ convergence for $1\leq k\leq\infty$, see Proposition 4.4 in [11].
As already mentioned, the risk measure VaR is well defined on $L^{0}$ and, in contrast to CVaR, is not a convex risk measure. It satisfies first order stochastic dominance, but fails to respect second order stochastic dominance. These features make optimization problems involving VaR particularly challenging.
The remainder of the paper is structured as follows. In Section 2, we analyze stochastic dominance relations within one-dimensional NMVM models. Section 3 is devoted to the study of frontier portfolios. In Section 4, we derive the corresponding capital asset pricing model.
2 High-order stochastic dominance
Stochastic dominance (SD) relations between random variables are characterized through pointwise comparisons of their distribution functions. This notion plays a central role in economics and finance, with applications in option pricing, portfolio insurance, and risk management. Therefore, determining necessary and sufficient conditions for SD between random variables is highly relevant. In this section, we focus on deriving such conditions for higher-order SD relations within the class of one-dimensional NMVM models.
Let $U_{k}=\{U:\mathbb{R}\rightarrow\mathbb{R}:(-1)^{j-1}U^{(j)}(x)\geq 0,\ \forall x\in\mathbb{R},\ 1\leq j\leq k\}$ denote the set of $k$-times differentiable utility functions whose derivatives alternate in sign, for each $k\geq 1$. We say that a scalar random variable $\eta_{1}$ stochastically dominates another scalar random variable $\eta_{2}$ in the $n$th order if $EU(\eta_{1})\geq EU(\eta_{2})$ for all $U\in U_{k}$. Throughout the paper, we write $\eta_{1}\succeq_{(k)}\eta_{2}$ to represent this SD relation.
This SD relation among random variables can also be characterized through their higher-order cumulative distribution functions. For an arbitrary real-valued random variable $\eta$, let $F^{(1)}_{\eta}(x)$ denote its right-continuous distribution function. Then, for each integer $k\geq 2$, define recursively
$F^{(k)}_{\eta}(x)=\int_{-\infty}^{x}F^{(k-1)}_{\eta}(s)\,ds,\;\;\forall x\in\mathbb{R}.$
We refer to $F^{(k)}_{\eta}(x)$ as the $k$th-order cumulative distribution function of $\eta$, and we will use the shorthand notation $k$-CDF for brevity. For any two random variables $\eta_{1}$ and $\eta_{2}$, the relation $\eta_{1}\succeq_{(k)}\eta_{2}$ can be equivalently expressed as
$F^{(k)}_{\eta_{1}}(x)\leq F^{(k)}_{\eta_{2}}(x),\;\;\forall x\in\mathbb{R}.$ (3)
If strict inequality in (3) holds for at least one point $x$, we write $\eta_{1}\succ_{(k)}\eta_{2}$. Evidently, condition (3) is intrinsically high dimensional, as the inequality must be verified for every real number $x$. Consequently, analyzing kSD relations among random variables is a nontrivial task. In this section, in order to obtain results on kSD relations for NMVM models, we verify condition (3) directly, after deriving and studying the properties of their $k$-CDFs for all positive integers $k$. For additional discussion of these kSD dominance relations, we refer to [3], [4], [25], [26] and the references therein.
2.1 High-order CDFs
The main objective of this subsection is to derive explicit formulas for high-order CDFs corresponding to normal, elliptical, and NMVM models. Our computations in this part are largely based on key characterizations of the kSD property established in [26]. We begin by recalling several results from [26] that will be crucial for our analysis, and we adopt the same notation as in that paper.
For each $k\in\mathbb{N}$ (where $\mathbb{N}$ denotes the set of positive integers throughout) and any random variable $\eta\in L^{k}$, define
[…]
where $x^{+}$ denotes the positive part of $x$. For any integer $k\geq 2$ and any pair $\eta_{1},\eta_{2}\in L^{k}$, Proposition 1 in [26] shows that kSD is equivalent to
$G_{\eta_{1}}^{(k-1)}(x)\leq G_{\eta_{2}}^{(k-1)}(x),\;\forall x\in\mathbb{R}.$ (4)
Furthermore, by Proposition 6 in [26], the function $G_{\eta}^{(k)}(x)$ is increasing and convex, satisfies $\lim_{x\rightarrow-\infty}G_{\eta}^{(k)}(x)=0$, and obeys $G_{\eta}^{(k)}(x)\geq x-E\eta$ for all $x\in\mathbb{R}$. The same proposition also states that $x-E\eta$ is a right asymptote of $G_{\eta}^{(k)}(x)$, in the sense that
$\lim_{x\rightarrow+\infty}\left(G_{\eta}^{(k)}(x)-(x-E\eta)\right)=0.$
In the remainder of this section, we explicitly compute the $k$-CDFs for several classes of random variables for arbitrary $k\in\mathbb{N}$. These expressions are naturally linked to the kSD property via (3).
Normal random variables: For a standard normal random variable $\eta\sim N(0,1)$, we write $F_{\eta}^{(k)}(x)$ as $\phi^{(k)}(x)$. In particular,
$\phi^{(k)}(x)=\frac{1}{(k-1)!}E[(x-N)^{+}]^{k-1},$ (5)
where $N$ denotes a standard normal random variable. We now establish the following elementary lemma.
Lemma 2.1.
For every integer $k\geq 2$, the following relation holds:
$\phi^{(k)}(x)=\frac{1}{k-1}\bigl[x\,\phi^{(k-1)}(x)+\phi^{(k-2)}(x)\bigr],$ (6)
where $\phi^{(0)}(x)=\phi(x)$ denotes the probability density function of a standard normal random variable, and $\phi^{(1)}(x)=\varphi(x)$ denotes its cumulative distribution function.
Proof.
We argue by induction. For $k=2$, we compute
$\begin{split}\phi^{(2)}(x)=&\int_{-\infty}^{x}\phi^{(1)}(s)\,ds=s\phi^{(1)}(s)\big|_{-\infty}^{x}-\int_{-\infty}^{x}s\phi^{(0)}(s)\,ds\\ =&\,x\phi^{(1)}(x)+\phi^{(0)}(x),\end{split}$
where we used that $\lim_{s\rightarrow-\infty}[s\phi^{(1)}(s)]=\lim_{s\rightarrow-\infty}\phi^{(1)}(s)/\frac{1}{s}=0$, which follows from L’Hôpital’s rule. Now suppose that (6) holds for $k$; we show that it then holds for $k+1$. To this end, we integrate both sides of (6) and afterwards substitute (6). This yields
$\begin{split}\phi^{(k+1)}(x)=&\int_{-\infty}^{x}\phi^{(k)}(s)\,ds=\frac{1}{k-1}\int_{-\infty}^{x}s\phi^{(k-1)}(s)\,ds+\frac{1}{k-1}\phi^{(k-1)}(x)\\ =&\frac{1}{k-1}\big[s\phi^{(k)}(s)\big]\big|_{-\infty}^{x}-\frac{1}{k-1}\phi^{(k+1)}(x)+\frac{1}{k-1}\phi^{(k-1)}(x)\\ =&\frac{1}{k-1}\big[x\phi^{(k)}(x)\big]-\frac{1}{k-1}\phi^{(k+1)}(x)+\frac{1}{k-1}\phi^{(k-1)}(x),\end{split}$ (7)
where we have used $\lim_{s\rightarrow-\infty}[s\phi^{(k)}(s)]=\lim_{s\rightarrow-\infty}\phi^{(k)}(s)/\frac{1}{s}=0$, which follows from repeated applications of L’Hôpital’s rule. Rearranging (7) gives
$\phi^{(k+1)}(x)=\frac{1}{k}\big[x\phi^{(k)}(x)+\phi^{(k-1)}(x)\big],$
which establishes the induction step and completes the proof. ∎
Remark 2.2.
From relation (6) we obtain
[…]
We see that the sequence $\phi^{(k)}(0)$ is monotone decreasing and converges to zero. Moreover, from (6) we deduce that $\phi^{(k)}(x)\geq\frac{x}{k-1}\phi^{(k-1)}(x)$, and hence, for $x\geq k-1$ it follows that $\phi^{(k)}(x)\geq\phi^{(k-1)}(x)$. Consequently, $\phi^{(k)}(x)$ and $\phi^{(k-1)}(x)$ must intersect at some point $x>0$.
Next, using (6), we obtain the following representation for $\phi^{(k)}(x)$.
Lemma 2.3.
For any $k\geq 2$,
$\phi^{(k)}(x)=p_{k-1}(x)\varphi(x)+q_{k-2}(x)\phi(x),$ (8)
where $p_{k-1}(x)$ is a polynomial of degree k1k-1 satisfying
$p_{j}(x)=\frac{x}{j}p_{j-1}(x)+\frac{1}{j}p_{j-2}(x),\;\;j\geq 2,\;\;\;p_{1}(x)=x,\;\;p_{0}(x)=1,$
and $q_{k-2}(x)$ is a polynomial of degree k2k-2 satisfying
$q_{i}(x)=\frac{x}{i+1}q_{i-1}(x)+\frac{1}{i+1}q_{i-2}(x),\;\;i\geq 2,\;\;\;q_{1}(x)=\frac{x}{2},\;\;q_{0}(x)=1.$
Proof.
For $k=2$, (6) immediately yields
$\phi^{(2)}(x)=x\varphi(x)+\phi(x)=p_{1}(x)\varphi(x)+q_{0}(x)\phi(x).$
For $k=3$, a direct computation gives
$\phi^{(3)}(x)=\frac{x^{2}+1}{2}\phi^{(2)}(x)+\frac{x}{2}\phi^{(0)}(x)=p_{2}(x)\varphi(x)+q_{1}(x)\phi(x),$
again by applying (6). Now assume that (8) holds for all integers $2,3,\dots,k$, and we show it holds for $k+1$. From (6),
$\begin{split}\phi^{(k+1)}(x)&=\frac{x}{k}\phi^{(k)}(x)+\frac{1}{k}\phi^{(k-1)}(x)\\ &=\frac{x}{k}[p_{k-1}(x)\varphi(x)+q_{k-2}(x)\phi(x)]+\frac{1}{k}[p_{k-2}(x)\varphi(x)+q_{k-3}(x)\phi(x)]\\ &=\bigg[\frac{x}{k}p_{k-1}(x)+\frac{1}{k}p_{k-2}(x)\bigg]\varphi(x)+\bigg[\frac{x}{k}q_{k-2}(x)+\frac{1}{k}q_{k-3}(x)\bigg]\phi(x)\\ &=p_{k}(x)\varphi(x)+q_{k-1}(x)\phi(x),\end{split}$
where we define
$p_{k}(x):=\frac{x}{k}p_{k-1}(x)+\frac{1}{k}p_{k-2}(x),\qquad q_{k-1}(x):=\frac{x}{k}q_{k-2}(x)+\frac{1}{k}q_{k-3}(x).$
It is evident that $p_{k}(x)$ is a polynomial of degree $k$ and $q_{k-1}(x)$ is a polynomial of degree k1k-1. ∎
Lemma 2.4.
The function $y(x)=\phi^{(k)}(x)$ satisfies
[…]
and the polynomial solution $y(x)=\sum_{j=0}^{+\infty}a_{j}x^{j}$ of (9) is determined by
$a_{j+3}=\frac{(k-1)-(j+1)}{(j+2)(j+3)}a_{j+1},\;\;j=0,1,\cdots,$
with initial coefficients
$a_{2}=\frac{k-1}{2}a_{0},\;\;a_{1}=\phi^{(k-1)}(0),\;\;a_{0}=\phi^{(k)}(0).$
Proof.
The relation (6) can be rewritten in the form
$(k-1)\phi^{(k)}(x)=x\phi^{(k-1)}(x)+\phi^{(k-2)}(x).$
Defining $y:=\phi^{(k)}(x)$, we have $y^{\prime}=\phi^{(k-1)}(x)$ and $y^{\prime\prime}=\phi^{(k-2)}(x)$, so the differential equation in (9) follows immediately. To determine its polynomial solution, we substitute $y(x)=\sum_{j=0}^{+\infty}a_{j}x^{j}$ into (9), which yields a polynomial identically equal to zero. Consequently, each coefficient of this resulting polynomial must vanish, leading to the stated recurrence for $a_{j}$. ∎
Remark 2.5.
Note that $a_{k+1}=0$, and thus $a_{k+2j+1}=0$ for every $j\geq 0$. Moreover,
$a_{k+2j}=(-1)^{j}\frac{3\cdot 5\cdots(2j-3)\cdot(2j-1)}{(k+2j)!}\frac{1}{\sqrt{2\pi}},$
and
$a_{0}=\phi^{(k)}(0),\;a_{1}=\phi^{(k-1)}(0),\;\cdots,\;a_{j}=\frac{1}{j!}\phi^{(k-j)}(0),\;\cdots,\;a_{k}=\frac{1}{k!}\phi^{(0)}(0)=\frac{1}{k!}\frac{1}{\sqrt{2\pi}}.$
Remark 2.6.
In the case that $\eta\sim N(u,\sigma^{2})$, we denote $F_{\eta}^{(k)}(x)$ by $\varphi^{(k)}(x;u,\sigma^{2})$. By using (5), we can easily obtain
$\varphi^{(k)}(x;u,\sigma^{2})=\sigma^{k-1}\phi^{(k)}(\frac{x-u}{\sigma}).$
By letting $y(x)=\varphi^{(k)}(x;u,\sigma^{2})$, one can easily show that it satisfies the following equation
$\sigma^{2}y^{{}^{\prime\prime}}+(x-u)y^{\prime}-(k-1)y=0,$
with the initial conditions
$y(0)=\sigma^{k-1}\phi^{(k)}(-\frac{u}{\sigma}),\;y^{\prime}(0)=\sigma^{k-2}\phi^{(k-1)}(-\frac{u}{\sigma}).$
Elliptical random variables: Suppose $\eta\sim u+\sigma\sqrt{Z}N(0,1)$, where $Z\in L^{k-1}$ is a positive random variable independent of $N(0,1)$. In this case, we write $F^{(k)}_{\eta}(x)$ as $\varphi_{e}^{(k)}(x;u,\sigma)$. By (5), we obtain
$\varphi_{e}^{(k)}(x;u,\sigma)=\sigma^{k-1}\int_{0}^{+\infty}z^{\frac{k-1}{2}}\phi^{(k)}\!\Big(\frac{x-u}{\sigma\sqrt{z}}\Big)f_{Z}(z)\,dz.$
Lemma 2.7.
For every $k\geq 2$, if $Z\in L^{k-1}$ then
$\begin{split}\frac{d\varphi_{e}^{(k)}(x;u,\sigma)}{du}=&-\sigma^{k-2}\int_{0}^{+\infty}z^{\frac{k-2}{2}}\phi^{(k-1)}\!\Big(\frac{x-u}{\sigma\sqrt{z}}\Big)f_{Z}(z)\,dz<0,\\[5.69046pt] \frac{d\varphi_{e}^{(k)}(x;u,\sigma)}{d\sigma}=&\sigma^{k-2}\int_{0}^{+\infty}z^{\frac{k-1}{2}}\phi^{(k-2)}\!\Big(\frac{x-u}{\sigma\sqrt{z}}\Big)f_{Z}(z)\,dz>0.\end{split}$ (10)
Proof.
The first identity in (10) follows immediately. For the second one, observe that
$\begin{split}\frac{d\varphi_{e}^{(k)}(x;u,\sigma)}{d\sigma}=&(k-1)\sigma^{k-2}\int_{0}^{+\infty}z^{\frac{k-1}{2}}\phi^{(k)}\!\Big(\frac{x-u}{\sigma\sqrt{z}}\Big)f_{Z}(z)\,dz\\ &+\sigma^{k-1}\int_{0}^{+\infty}z^{\frac{k-1}{2}}\Big(-\frac{x-u}{\sigma^{2}\sqrt{z}}\Big)\phi^{(k-1)}\!\Big(\frac{x-u}{\sigma\sqrt{z}}\Big)f_{Z}(z)\,dz\\ =&\sigma^{k-2}\int_{0}^{+\infty}z^{\frac{k-1}{2}}\Big[(k-1)\phi^{(k)}\!\Big(\frac{x-u}{\sigma\sqrt{z}}\Big)-\Big(\frac{x-u}{\sigma\sqrt{z}}\Big)\phi^{(k-1)}\!\Big(\frac{x-u}{\sigma\sqrt{z}}\Big)\Big]f_{Z}(z)\,dz.\end{split}$ (11)
Using (6), we obtain
$(k-1)\phi^{(k)}\!\Big(\frac{x-u}{\sigma\sqrt{z}}\Big)-\frac{x-u}{\sigma\sqrt{z}}\phi^{(k-1)}\!\Big(\frac{x-u}{\sigma\sqrt{z}}\Big)=\phi^{(k-2)}\!\Big(\frac{x-u}{\sigma\sqrt{z}}\Big).$ (12)
Substituting (12) into (11) yields the second identity in the proposition. ∎
NMVM random variables: When $\eta\sim a+bZ+\sigma\sqrt{Z}N(0,1)$, we denote $F^{(k)}_{\eta}(x)$ by $\varphi_{h}^{(k)}(x;a,b,\sigma)$ (using the initial letter of “hyperbolic” distributions). From (5), it follows that
$\varphi_{h}^{(k)}(x;a,b,\sigma)=\sigma^{k-1}\int^{+\infty}_{0}z^{\frac{k-1}{2}}\phi^{(k)}\!\left(\frac{x-a-bz}{\sigma\sqrt{z}}\right)f_{Z}(z)\,dz.$ (13)
By straightforward differentiation and by invoking (6), we arrive at the following result.
Lemma 2.8.
For any integer $k\geq 2$, if $Z\in L^{k-1}$, then
[…]
Proof.
The first two identities in (14) are obtained by differentiating (13) with respect to $a$ and $b$, respectively. The third identity follows by differentiating (13) with respect to $\sigma$ and then applying (6) in the same way as in the proof of Proposition 2.7. ∎
Remark 2.9.
Lemma 2.8 implies that for any two random variables $\eta_{1}\sim a_{1}+b_{1}Z+\sigma_{1}\sqrt{Z}N(0,1)$ and $\eta_{2}\sim a_{2}+b_{2}Z+\sigma_{2}\sqrt{Z}N(0,1)$, the condition $a_{1}\geq a_{2},\ b_{1}\geq b_{2},\ \sigma_{2}\geq\sigma_{1}>0$ yields $\eta_{1}\succeq_{(k)}\eta_{2}$. Moreover, if at least one of these inequalities is strict, then $\eta_{1}\succ_{(k)}\eta_{2}$.
2.2 High-order SD
As mentioned earlier, the condition (3), which must be verified for all real numbers $x$, indicates that SD is an infinite-dimensional problem. Consequently, it is difficult to derive necessary and sufficient conditions for SD in general. For certain special classes of random variables, however, various characterizations of second-order stochastic dominance are well established in the literature. For instance, if $N_{1}\sim N(u_{1},\sigma_{1})$ and $N_{2}\sim N(u_{2},\sigma_{2})$, then $N_{1}\succeq_{(2)}N_{2}$ if and only if $u_{1}\geq u_{2}$ and $\sigma_{1}\leq\sigma_{2}$; see, for example, Proposition 2.59 in [13] and Theorem 6.2 in [20]. For general random variables such convenient necessary and sufficient criteria for SD are, however, hard to obtain. In this section we provide some necessary and some sufficient conditions for kSD in the setting of NMVM models.
By definition, the $n_{1}$-th order SD relation is weaker than the $n_{2}$-th order SD relation whenever $n_{1}\geq n_{2}$, since $U_{n_{1}}\subset U_{n_{2}}$. Thus it is not obvious that, for normal random variables, $N_{1}\succeq_{(k)}N_{2}$ would still imply $u_{1}\geq u_{2}$ and $\sigma_{1}\leq\sigma_{2}$ for $k\geq 3$. In this section, as a consequence of a more general result for NVMM, we prove that, in fact, $N_{1}\succeq_{(k+1)}N_{2}$ is equivalent to $u_{1}\geq u_{2}$ and $\sigma_{1}\leq\sigma_{2}$ for every $k\in\mathbb{N}$; see Proposition 2.10 below. Our argument provides an alternative proof of the classical second-order stochastic dominance characterization for normal random variables (cf. Theorem 6.2 in [20] and Theorem 3.1 in [5]), and at the same time extends this type of characterization to kSD relations (in particular for $k\geq 3$) for elliptical random variables.
Before turning to the main calculations, we recall a pertinent result from [26]. With $G_{\eta}^{(k)}(x)$ defined as in Subsection 2.1, the value of $G_{\eta}^{(k)}(x)$ at $x=E\eta$ is called the central semideviation of $\eta$ and is denoted by $\bar{\delta}_{\eta}^{(k)}=G_{\eta}^{(k)}(E\eta)$ (we follow the notation of [26]). Explicitly,
$\bar{\delta}_{\eta}^{(k)}=\|(E\eta-\eta)^{+}\|_{k}=\bigl[E\bigl((E\eta-\eta)^{+}\bigr)^{k}\bigr]^{\frac{1}{k}}.$
Proposition 4 in [26] states that $\bar{\delta}_{\eta}^{(k)}$ is convex, i.e.,
$\bar{\delta}_{t\eta_{1}+(1-t)\eta_{2}}^{(k)}\leq t\bar{\delta}_{\eta_{1}}^{(k)}+(1-t)\bar{\delta}_{\eta_{2}}^{(k)}$
for every $t\in[0,1]$ and all $\eta_{1},\eta_{2}\in L^{k}$. Furthermore, Corollary 2 of the same paper shows that if $\eta_{1}\succeq_{(k+1)}\eta_{2}$, then
$E\eta_{1}-\bar{\delta}_{\eta_{1}}^{(n)}\geq E\eta_{2}-\bar{\delta}_{\eta_{2}}^{(n)},\;\;\;E\eta_{1}\geq E\eta_{2},$ (15)
for all $n\geq k$, provided that $\eta_{1},\eta_{2}\in L^{n}$. This relation (15) will play a crucial role in the analysis carried out in this subsection.
As a direct application of this result, we first establish the following proposition.
Proposition 2.10.
Let $\eta_{1}\sim u_{1}+\sigma_{1}ZN(0,1)$ and $\eta_{2}\sim u_{2}+\sigma_{2}ZN(0,1)$ be two Elliptical random variables with $u_{1},u_{2}\in\mathbb{R}$, $\sigma_{1}>0$, $\sigma_{2}>0$, and $Z$ a positive random variable satisfying $EZ^{k}<\infty$ for all positive integers $k$. Then, for every $k\geq 2$, we have $\eta_{1}\succeq_{(k)}\eta_{2}$ if and only if $u_{1}\geq u_{2}$ and $\sigma_{1}\leq\sigma_{2}$.
Proof.
[…]
Next, we examine kSD relations within the family of one–dimensional NMVM models. We begin by setting up some notation. Let $g(\cdot)$ be an arbitrary Borel measurable function from $(0,+\infty)$ to $(0,+\infty)$, and define
$Z_{g}:=g(Z).$
For any real numbers $a_{1},b_{1},a_{2},b_{2}$, set $\bar{X}_{g}=a_{1}+b_{1}Z_{g}+\sqrt{Z}\,N$ and $\bar{Y}_{g}=a_{2}+b_{2}Z_{g}+\sqrt{Z}\,N$. We first establish the following lemma.
Lemma 2.11.
Assume that the mixing variable $Z$ satisfies $Z\in L^{k}$ for every positive integer $k$. Then, for any bounded Borel function $g$ and any real numbers $a_{1},b_{1},a_{2},b_{2}$, we have
$\lim_{k\to+\infty}\frac{\|(\bar{X}_{g})^{+}\|_{k}}{\|(\bar{Y}_{g})^{+}\|_{k}}=1.$ (17)
Proof.
Note that
$\bar{X}_{g}-\bar{Y}_{g}=a_{1}-a_{2}+(b_{1}-b_{2})Z_{g}$
is a bounded random variable. Using $(a+b)^{+}\leq a^{+}+b^{+}$ together with the triangle inequality for norms, we obtain
$\|(\bar{X}_{g})^{+}\|_{k}=\|(\bar{X}_{g}-\bar{Y}_{g}+\bar{Y}_{g})^{+}\|_{k}\leq\|(\bar{X}_{g}-\bar{Y}_{g})^{+}\|_{k}+\|(\bar{Y}_{g})^{+}\|_{k}.$
Hence,
$\frac{\|(\bar{X}_{g})^{+}\|_{k}}{\|(\bar{Y}_{g})^{+}\|_{k}}\leq 1+\frac{\|(\bar{X}_{g}-\bar{Y}_{g})^{+}\|_{k}}{\|(\bar{Y}_{g})^{+}\|_{k}}.$ (18)
Since $(\bar{X}_{g}-\bar{Y}_{g})^{+}$ is bounded, we have $\sup_{k\geq 1}\|(\bar{X}_{g}-\bar{Y}_{g})^{+}\|_{k}<\infty$, while $(\bar{Y}_{g})^{+}$ is unbounded, so $\|(\bar{Y}_{g})^{+}\|_{k}\to\infty$ as $k\to\infty$. Therefore, from (18) it follows that
$\overline{\lim}_{k\to\infty}\frac{\|(\bar{X}_{g})^{+}\|_{k}}{\|(\bar{Y}_{g})^{+}\|_{k}}\leq 1.$ (19)
Applying the same reasoning in the reverse direction, we get
$\|(\bar{Y}_{g})^{+}\|_{k}=\|(\bar{Y}_{g}-\bar{X}_{g}+\bar{X}_{g})^{+}\|_{k}\leq\|(\bar{X}_{g}-\bar{Y}_{g})^{+}\|_{k}+\|(\bar{X}_{g})^{+}\|_{k}.$
Thus,
$\frac{\|(\bar{X}_{g})^{+}\|_{k}}{\|(\bar{Y}_{g})^{+}\|_{k}}\geq\frac{\|(\bar{X}_{g})^{+}\|_{k}}{\|(\bar{X}_{g}-\bar{Y}_{g})^{+}\|_{k}+\|(\bar{X}_{g})^{+}\|_{k}}=\frac{1}{\|(\bar{X}_{g}-\bar{Y}_{g})^{+}\|_{k}/\|(\bar{X}_{g})^{+}\|_{k}+1}.$ (20)
Since $(\bar{X}_{g})^{+}$ is unbounded, we have $\|(\bar{X}_{g})^{+}\|_{k}\to\infty$ as $k\to\infty$, and therefore
$\frac{\|(\bar{X}_{g}-\bar{Y}_{g})^{+}\|_{k}}{\|(\bar{X}_{g})^{+}\|_{k}}\to 0\quad\text{as }k\to\infty.$
Consequently, from (20) we obtain
$\underline{\lim}_{k\to\infty}\frac{\|(\bar{X}_{g})^{+}\|_{k}}{\|(\bar{Y}_{g})^{+}\|_{k}}\geq 1.$ (21)
Combining (19) and (21) yields (17). ∎
For any bounded, positive Borel function $g$ and real numbers $a_{1},a_{2},b_{1},b_{2}$ with $c_{1}>0$ and $c_{2}>0$, define
$\tilde{X}_{g}=a_{1}+b_{1}Z_{g}+c_{1}\sqrt{Z}\,N,\qquad\tilde{Y}_{g}=a_{2}+b_{2}Z_{g}+c_{2}\sqrt{Z}\,N.$
The next proposition is the main result of this section.
Proposition 2.12.
Consider model (1) and suppose that $Z\in L^{k}$ for every positive integer $k$. Then, for any positive bounded Borel function $g$ and each $k\in\mathbb{N}$, the relation $\tilde{X}_{g}\succeq_{(k+1)}\tilde{Y}_{g}$ implies
$a_{1}+b_{1}EZ_{g}\;\geq\;a_{2}+b_{2}EZ_{g}\quad\text{and}\quad c_{1}\leq c_{2}.$
Proof.
Since
$E\tilde{X}_{g}=a_{1}+b_{1}EZ_{g},\qquad E\tilde{Y}_{g}=a_{2}+b_{2}EZ_{g},$
the inequality $a_{1}+b_{1}EZ_{g}\geq a_{2}+b_{2}EZ_{g}$ follows directly from Theorem 1 in [26]. To establish $c_{1}\leq c_{2}$, we invoke Corollary 2 of the same reference [26].
First, note that for each integer $j>k$ the central semi-deviation of order $j$ is given by
$\bar{\delta}_{\tilde{X}_{g}}^{(j)}=c_{1}\left\|\Big(\frac{b_{1}}{c_{1}}EZ_{g}-\frac{b_{1}}{c_{1}}Z_{g}+ZN\Big)^{+}\right\|_{j},\qquad\bar{\delta}_{\tilde{Y}_{g}}^{(j)}=c_{2}\left\|\Big(\frac{b_{2}}{c_{2}}EZ_{g}-\frac{b_{2}}{c_{2}}Z_{g}+ZN\Big)^{+}\right\|_{j}.$
Set
$D_{j}:=\left\|\Big(\frac{b_{1}}{c_{1}}EZ_{g}-\frac{b_{1}}{c_{1}}Z_{g}+ZN\Big)^{+}\right\|_{j},\qquad E_{j}:=\left\|\Big(\frac{b_{2}}{c_{2}}EZ_{g}-\frac{b_{2}}{c_{2}}Z_{g}+ZN\Big)^{+}\right\|_{j}.$
By Lemma 2.11 we have $\lim_{j\to\infty}D_{j}/E_{j}=1$. Moreover, since the random variables
$\Big(\frac{b_{1}}{c_{1}}EZ_{g}-\frac{b_{1}}{c_{1}}Z_{g}+ZN\Big)^{+},\quad\Big(\frac{b_{2}}{c_{2}}EZ_{g}-\frac{b_{2}}{c_{2}}Z_{g}+ZN\Big)^{+}$
are unbounded, it follows that $\lim_{j\to\infty}D_{j}=+\infty$ and $\lim_{j\to\infty}E_{j}=+\infty$.
Corollary 2 in [26] then yields
$a_{1}+b_{1}EZ_{g}-c_{1}D_{j}\;\geq\;a_{2}+b_{2}EZ_{g}-c_{2}E_{j},$ (22)
for all $j\geq k$. Dividing both sides of (22) by $E_{j}$ and letting $j\to\infty$, we obtain $c_{1}\geq-c_{2}$, which is equivalent to $c_{1}\leq c_{2}$. ∎
Now, for any positive Borel function $g$, any real numbers $b_{1},b_{2}$, and any $c_{1}>0,c_{2}>0$, define
$\hat{X}_{g}=b_{1}Z_{g}+c_{1}\sqrt{Z}N,\qquad\hat{Y}_{g}=b_{2}Z_{g}+c_{2}\sqrt{Z}N.$
We obtain the following corollary.
Corollary 2.13.
Assume that $Z\in L^{k}$ for all positive integers $k$. Then, for any positive, bounded Borel function $g$ and each $k\in\mathbb{N}$,
$\hat{X}_{g}\succeq_{(k+1)}\hat{Y}_{g}\quad\Longleftrightarrow\quad b_{1}\geq b_{2},\;\;c_{1}\leq c_{2}.$
Proof.
If $\hat{X}_{g}\succeq_{(k+1)}\hat{Y}_{g}$, then Proposition 2.12 gives $b_{1}EZ_{g}\geq b_{2}EZ_{g}$ and $c_{1}\leq c_{2}$. Since $EZ_{g}>0$, this is equivalent to $b_{1}\geq b_{2}$ together with $c_{1}\leq c_{2}$.
For the converse, consider
$Y_{g}:=bZ_{g}+\sigma\sqrt{Z}N,\qquad b\in\mathbb{R},\;\sigma>0.$
Then
$\varphi_{Y_{g}}^{(k)}(x;b,\sigma)=\sigma^{k-1}\int_{0}^{+\infty}z^{\frac{k-1}{2}}\phi^{(k)}\!\left(\frac{x-bz_{g}}{\sigma\sqrt{z}}\right)f_{Z}(z)\,dz,$
where $z_{g}=g(z)$. Proceeding as in the proof of Proposition 2.8, we obtain
$\begin{split}\frac{d}{db}\varphi_{Y_{g}}^{(k)}(x;b,\sigma)&=-\,\sigma^{k-2}\int_{0}^{+\infty}z^{\frac{k-2}{2}}z_{g}\,\phi^{(k-1)}\!\left(\frac{x-bz_{g}}{\sigma\sqrt{z}}\right)f_{Z}(z)\,dz<0,\\[4.0pt] \frac{d}{d\sigma}\varphi_{Y_{g}}^{(k)}(x;b,\sigma)&=\sigma^{k-2}\int_{0}^{+\infty}z^{\frac{k-1}{2}}\phi^{(k-2)}\!\left(\frac{x-bz_{g}}{\sigma\sqrt{z}}\right)f_{Z}(z)\,dz>0.\end{split}$
These monotonicity properties imply that $b_{1}\geq b_{2}$ and $c_{1}\leq c_{2}$ lead to
$\hat{X}_{g}\succeq_{(k+1)}\hat{Y}_{g}.$
3 Frontier portfolios
In this section, we investigate solutions to problem (2). For ease of exposition, we first fix a positive integer $k\geq 1$ and consider risk measures $\rho:L^{k}\to\mathbb{R}\cup\{\infty\}$ associated with problem (2). Writing $dom(\rho)=\{\eta\in L^{k}:\rho(\eta)<\infty\}$ for the effective domain of $\rho$, we say that $\rho$ is proper if $\rho(\eta)>$-$\infty$ for all $\eta\in L^{k}$ and $dom(\rho)\neq\emptyset$.
A risk measure on $L^{k}$ is called a convex risk measure if it is a proper monetary risk measure (that is, it is monotone and cash-invariant) and is convex, meaning
$\rho(\lambda\eta_{1}+(1-\lambda)\eta_{2})\leq\lambda\rho(\eta_{1})+(1-\lambda)\rho(\eta_{2}),\quad\forall\lambda\in[0,1],\ \forall\eta_{1},\eta_{2}\in L^{k}.$
A convex risk measure is said to be coherent if it is positively homogeneous, i.e., $\rho(\lambda\eta)=\lambda\rho(\eta)$ for all $\lambda>0$. By Corollary 2.3 of [17], any finite convex risk measure on $L^{k}$ is continuous on $L^{k}$. In particular, the risk measure $CVaR$ is a coherent risk measure on $L^{k}$ and is therefore continuous on $L^{k}$.
Another key property of $\rho$ for our analysis in this section is its compatibility with second order stochastic dominance. We say that $\rho$ is consistent with second order stochastic dominance (SSD-consistent) if, for any $\eta_{1},\eta_{2}\in L^{k}$, the relation $\eta_{1}\succeq_{(2)}\eta_{2}$ implies $\rho(\eta_{1})\geq\rho(\eta_{2})$. This SSD-consistency of $\rho$ will be needed in the proof of Theorem 3.4 below. We call $\rho$ law-invariant if $\rho(\eta_{1})=\rho(\eta_{2})$ whenever $\eta_{1}\overset{d}{=}\eta_{2}$. Every finite-valued, law-invariant, convex risk measure is SSD-consistent; see [23] and the references therein.
Remark 3.1.
We point out that a solution to problem (2) always exists when the risk measure $\rho$ is finite-valued, law-invariant, and convex. To verify this, first note that the mapping $\omega\mapsto\rho(-\omega^{T}\xi)$ is continuous whenever $\xi\in L_{d}^{k}$, since $\rho$ is continuous on $L^{k}$ as discussed above. Moreover, for any integrable random variable $\eta$, one has $E\eta\succeq_{(2)}\eta$, and thus $\rho(-\eta)\geq\rho(-E\eta)$; see Appendix B of [14] and Corollary 5.1 of [12] for analogous arguments. Consequently, on the set $\{\omega\in\mathbb{R}^{d}:E(-\omega^{T}X)=r,\ \omega^{T}e=1\}$ we have $r\succeq_{(2)}\omega^{T}X$, which yields
$\rho(-\omega^{T}X)\geq\rho(r)=r+\rho(0).$
Before presenting the main result of this section, we briefly recall the classical Markowitz mean–variance portfolio optimization problem. In the Markowitz framework, portfolio risk is quantified by the variance. The associated optimization problem is
[…]
A closed-form solution to (23) is well known and can be found in any standard textbook on the capital asset pricing model; see, for instance, page 64 of [16]. We record the solution here because it will be used in our main results. Let $\mu=EX$ be the mean vector and let $V=Cov(X)$ be the covariance matrix of $X$. For each $r\in\mathbb{R}$, the optimizer of (23) is
$\omega^{\star}_{r}=\omega^{\star}_{r}(\mu,V)=\frac{1}{d_{4}}\big[d_{2}(V^{-1}\mathbf{1})-d_{1}(V^{-1}\mu)\big]+\frac{r}{d_{4}}\big[d_{3}(V^{-1}\mu)-d_{1}(V^{-1}\mathbf{1})\big],$ (24)
where
$d_{1}=\mathbf{1}^{T}V^{-1}\mu,\quad d_{2}=\mu^{T}V^{-1}\mu,\quad d_{3}=\mathbf{1}^{T}V^{-1}\mathbf{1},\quad d_{4}=d_{2}d_{3}-d_{1}^{2}.$
Remark 3.2.
Since $V^{-1}$ is positive definite, we obtain $\big(d_{1}\mu-d_{2}\mathbf{1}\big)^{T}V^{-1}\big(d_{1}\mu-d_{2}\mathbf{1}\big)>0$ and $d_{2}>0$ (here we assume $\mu=EX\neq 0$). Moreover,
$\big(d_{1}\mu-d_{2}\mathbf{1}\big)^{T}V^{-1}\big(d_{1}\mu-d_{2}\mathbf{1}\big)=d_{2}(d_{2}d_{3}-d_{1}^{2})=d_{2}d_{4}>0,$
which implies $d_{4}>0$ in (24).
In this section, we demonstrate that the optimal portfolio for our mean–risk portfolio optimization problem (2) can also be written in a form analogous to (24). To this end, we begin by introducing some notation. For any random vector $\theta$ with mean vector $\mu_{\theta}=E\theta$ and covariance matrix $\Sigma_{\theta}=\mathrm{Cov}(\theta)$, define
$\omega^{\star}_{\theta}=\omega^{\star}_{\theta}(\mu_{\theta},\Sigma_{\theta})=\frac{1}{d_{4\theta}}\bigl[d_{2\theta}(\Sigma_{\theta}^{-1}\mathbf{1})-d_{1\theta}(\Sigma_{\theta}^{-1}\mu_{\theta})\bigr]+\frac{r}{d_{4\theta}}\bigl[d_{3\theta}(\Sigma_{\theta}^{-1}\mu_{\theta})-d_{1\theta}(\Sigma_{\theta}^{-1}\mathbf{1})\bigr],$ (25)
where
$d_{1\theta}=\mathbf{1}^{T}\Sigma_{\theta}^{-1}\mu_{\theta},\quad d_{2\theta}=\mu_{\theta}^{T}\Sigma_{\theta}^{-1}\mu_{\theta},\quad d_{3\theta}=\mathbf{1}^{T}\Sigma_{\theta}^{-1}\mathbf{1},\quad d_{4\theta}=d_{2\theta}d_{3\theta}-d_{1\theta}^{2}.$
Remark 3.3.
The right-hand side of (25) depends only on the mean vector $\mu_{\theta}$ and covariance matrix $\Sigma_{\theta}$ of the random vector $\theta$. Consequently, $\omega^{\star}_{\theta}=\omega^{\star}_{\eta}$ whenever $\theta$ and $\eta$ share the same mean vectors and covariance matrices. We use the notation $\omega^{\star}_{\theta}$ to emphasize that this portfolio is the efficient frontier portfolio, in the mean–variance sense, corresponding to the return vector $\theta$. Note that the formula (25) is well defined whenever $\mu_{\theta}\neq 0$ and $\Sigma_{\theta}$ are positive definite (i.e., all eigenvalues are strictly positive), as discussed in Remark 3.2.
The next result, Theorem 3.4, establishes that for problem (2), frontier portfolios under any real-valued, law-invariant, convex risk measure always exist and can be written in the form (25) when the return vectors are specified by (1).
Theorem 3.4.
Fix an integer $k\geq 1$. Let $\rho$ be a finite-valued, law-invariant, convex risk measure on $L^{k}$. Suppose the return vector is given by (1) with $Z\in L^{k}$. Then, for each fixed real number $r$, the solution to problem (2) is given by $\omega^{\star}_{\theta}$ as in (25) with $\mu_{\theta}=\mu+\gamma EZ$ and $\Sigma_{\theta}=\Sigma$.
Proof.
Define $D_{r}=\{\omega\in\mathbb{R}^{d}:\omega^{T}(\mu+\gamma EZ)=r,\ \omega^{T}\mathbf{1}=1\}$. On $D_{r}$, the minimizer of the quadratic form $\omega^{T}\Sigma\omega$ is $\bar{\omega}:=\omega_{\theta}(\mu+\gamma EZ,\Sigma)$, as in (25). By Corollary 3.13, for every $\omega\in D_{r}$ we have $\bar{\omega}^{T}X\succ_{(2)}\omega^{T}X$. The assumptions on $\rho$ in the theorem ensure that $\rho$ is SSD-consistent, and therefore $\rho(-\bar{\omega}^{T}X)\leq\rho(-\omega^{T}X)$ for all $\omega\in D_{r}$. This establishes the claim. ∎∎
Remark 3.5.
Observe that $\mu_{X}=EX=\mu+\gamma EZ=\mu_{\theta}$. However, the covariance matrix $\mathrm{Cov}(X)=\gamma\gamma^{T}\mathrm{Var}(Z)+A^{T}AEZ$ differs from $\Sigma_{\theta}=A^{T}A$. Hence, the optimal solution of (23) is not the same as the optimal solution of (2). They are nevertheless closely related in that they differ only in the covariance matrices $\Sigma_{X}$ and $\Sigma_{\theta}$.
Remark 3.6.
The content of Theorem 3.4 can be summarized as follows: the mean–risk efficient frontier portfolios associated with any real-valued, law-invariant, convex risk measure $\rho$ can be obtained by solving a Markowitz mean–variance portfolio optimization problem with a suitably modified return vector $\theta$, as specified in Theorem 3.4.
The risk measure $CVaR_{\alpha}$, for each fixed $\alpha\in(0,1)$, is finite-valued, law-invariant, and convex. Consequently, the conclusion of Theorem 3.4 above applies to this risk measure as well. We record this in the following corollary.
Corollary 3.7.
Let $X$ be a return vector as in (1) with $Z\in L^{k}$ for some positive integer $k$. Then, for any fixed $\alpha\in(0,1)$, the closed-form solution to the optimization problem
[…]
is given by $\omega_{\theta}^{\star}$ in (25), where $\mu_{\theta}=EY=\mu+\gamma EZ$ and $\Sigma_{\theta}=A^{T}A$.
Remark 3.8.
Theorem 3.4 above can be used to streamline the computation of optimal $CVaR_{\alpha}$ values on certain portfolio domains when returns are normally distributed—a setting analyzed in Theorem 2 of [29] (see also [30]).
To see this, first note that if $Z=1$ in (1), then $X$ becomes a Normal random vector. For notational simplicity in what follows, set $\mu=0$ and write $X\sim N_{d}(\gamma,\Sigma)$. With all other model parameters fixed, $\gamma$ and $\Sigma$ are then fixed, and expression (25) depends only on the target expected return $r$. Thus, for convenience, denote by $\omega_{r}:=\omega_{\theta}^{\star}$ (for $\theta\overset{d}{=}N_{d}(\gamma,\Sigma)$ in (25)). We can rewrite $\omega_{r}$ as
$\omega_{r}=k_{1}(r)\Sigma^{-1}e+k_{2}(r)\Sigma^{-1}\gamma,$
where
$k_{1}(r)=\frac{d_{2\theta}}{d_{4\theta}}-r\frac{d_{1\theta}}{d_{4\theta}},\quad k_{2}(r)=r\frac{d_{3\theta}}{d_{4\theta}}-\frac{d_{1\theta}}{d_{4\theta}}.$
With this notation we have $\omega_{r}^{T}N_{d}\sim N(\omega_{r}^{T}\gamma,\sigma_{r}^{2})$, where
$\sigma_{r}^{2}:=\omega_{r}^{T}\Sigma\omega_{r}=k_{1}^{2}(r)e^{T}\Sigma^{-1}e+2k_{1}(r)k_{2}(r)e^{T}\Sigma^{-1}\gamma+k_{2}^{2}(r)\gamma^{T}\Sigma^{-1}\gamma.$ (27)
From [28] (see also equation (2) in [18]), for a Normal random variable $H\sim N(\delta,\sigma^{2})$ we have
$CVaR_{\alpha}(H)=\delta+\Bigg[\frac{1}{\sigma}\,\frac{\phi\big(\frac{z_{\alpha}-\delta}{\sigma}\big)}{1-\varphi\big(\frac{z_{\alpha}-\delta}{\sigma}\big)}\Bigg]\sigma^{2},$ (28)
where $z_{\alpha}$ is the $\alpha$–quantile of the standard Normal distribution $N(0,1)$. For any return level $r$, define
$h_{\alpha}(r):=r+\Bigg[\frac{1}{\sigma_{r}}\,\frac{\phi\big(\frac{z_{\alpha}-r}{\sigma_{r}}\big)}{1-\varphi\big(\frac{z_{\alpha}-r}{\sigma_{r}}\big)}\Bigg]\sigma_{r}^{2},$
with $\sigma_{r}$ as in (27).
Now consider the domain
$D=\{\omega:\omega^{T}\gamma\geq\bar{r}\},$
for some fixed real number $\bar{r}$. Let $D_{\gamma}:=\{\omega^{T}\gamma:\omega\in D\}$. Note that $D$ has the property that for any $r_{0}\in D_{\gamma}$, every portfolio $\omega$ satisfying $\omega^{T}\gamma=r_{0}$ belongs to $D$, i.e., $\omega\in D$. On this domain we obtain
$\min_{\omega\in D}CVaR_{\alpha}(\omega^{T}\theta)=\min_{r\in D_{\gamma}}h_{\alpha}(r),$ (29)
where $\theta\overset{d}{=}N_{d}(\gamma,\Sigma)$. The right-hand side of (29) is simply the minimization of a real-valued function over a subset of the real line, which is significantly simpler than minimizing $F_{\alpha}(\omega,a)$ in (4) of [29], as done in Theorem 2 of that paper.
The proof of Theorem 3.4 requires some preliminary results. We begin by presenting a property of the second-order CDF function $\phi^{(2)}(x)$ associated with the standard Normal random variable, formulated as a lemma. This statement corresponds to part P8 of Proposition 2 in [25]. The lemma is crucial for establishing Lemma 3.10 below.
Lemma 3.9.
For any fixed real number $x_{0}$ we have
  • i)
    $\phi^{(2)}(x)-\phi^{(2)}(x_{0})\leq x-x_{0}$ for all $x\geq x_{0}$.
  • ii)
    $\phi^{(2)}(x)-\phi^{(2)}(x_{0})\geq x-x_{0}$ for all $x\leq x_{0}$.
Lemma 3.10.
Let $Z$ be any positive bounded random variable, i.e., $m_{1}\leq Z\leq m_{2}$ for some finite numbers $0<m_{1}<m_{2}$. Denote $m=EZ$. Consider the one dimensional NMVM models $\eta_{1}=a_{1}+b_{1}Z+c_{1}\sqrt{Z}N(0,1)$ and $\eta_{2}=a_{2}+b_{2}Z+c_{2}\sqrt{Z}N(0,1)$, where $a_{1},a_{2},b_{1},b_{2},$ are any real numbers and $c_{1}>0,\;c_{2}>0$. If
$a_{1}+b_{1}EZ=a_{2}+b_{2}EZ,\;\;c_{1}<c_{2},$ (30)
then $\eta_{1}\succ_{(2)}\eta_{2}$.
Proof.
We need to show $I_{1}(x)=:E(x-\eta_{1})^{+}\leq I_{2}(x)=:E(x-\eta_{2})^{+}$ for any real number $x$ with strict inequality for some $x$. From (13) we have
$I_{1}(x)=\int_{\alpha}^{\beta}c_{1}\sqrt{z}\phi^{(2)}(\frac{x-a_{1}-b_{1}z}{c_{1}\sqrt{z}})f(z)dz,\;\;I_{2}(x)=\int_{\alpha}^{\beta}c_{2}\sqrt{z}\phi^{(2)}(\frac{x-a_{2}-b_{2}z}{c_{2}\sqrt{z}})f(z)dz.$
Further we have
[…]
Step 1: We first evaluate $J_{2}(x)$. We define the function $Q_{z}(c)=:c\sqrt{z}\phi^{(2)}(\frac{x-a_{1}-b_{1}z}{c\sqrt{z}})$ for each fixed $z,x,a_{1},b_{1}$. We have
$Q_{z}^{\prime}(c)=\sqrt{z}[\phi^{(2)}(\frac{x-a_{1}-b_{1}z}{c\sqrt{z}})-\frac{x-a_{1}-b_{1}z}{c\sqrt{z}}\phi^{(1)}(\frac{x-a_{1}-b_{1}z}{c\sqrt{z}})]=\sqrt{z}\phi^{(0)}(\frac{x-a_{1}-b_{1}z}{c\sqrt{z}}).$
Here the second equality follows from the relation $\phi^{(2)}(x)-x\phi(x)=\phi^{(0)}(x)$ due to (6). By the middle value Theorem, we have $Q_{z}(c_{2})-Q_{z}(c_{1})=\sqrt{z}\phi^{(0)}(\frac{x-a_{1}-b_{1}z}{\bar{c}\sqrt{z}})$ for some $c_{1}<\bar{c}<c_{2}$. Now if $x-a_{1}-b_{1}z\geq 0$, then $\frac{x-a_{1}-b_{1}z}{\bar{c}\sqrt{z}}\leq\frac{x-a_{1}-b_{1}z}{c_{1}\sqrt{z}}$ and therefore $\phi^{(0)}(\frac{x-a_{1}-b_{1}z}{\bar{c}\sqrt{z}})\geq\phi^{(0)}(\frac{x-a_{1}-b_{1}z}{c_{1}\sqrt{z}})$ (recall here that $\phi^{(0)}(x)=\frac{1}{\sqrt{2\pi}}e^{-\frac{x^{2}}{2}}$). If $x-a_{1}-b_{1}z<0$, then $|\frac{x-a_{1}-b_{1}z}{\bar{c}\sqrt{z}}|\leq|\frac{x-a_{1}-b_{1}z}{c_{1}\sqrt{z}}|$ and therefore we still have $\phi^{(0)}(\frac{x-a_{1}-b_{1}z}{\bar{c}\sqrt{z}})\geq\phi^{(0)}(\frac{x-a_{1}-b_{1}z}{c_{1}\sqrt{z}})$. From these we conclude that
$J_{2}(x)=\int_{\alpha}^{\beta}[Q_{z}(c_{2})-Q_{z}(c_{1})]f(z)dz\geq(c_{2}-c_{1})\int_{\alpha}^{\beta}\sqrt{z}\phi^{(0)}(\frac{x-a_{1}-b_{1}z}{c_{1}\sqrt{z}})f(z)dz$
Step 2: In this step we show $J_{1}(x)\geq 0$ for any real number $x$. Define $\bar{\eta}_{1}=a_{1}+b_{1}Z+c_{2}\sqrt{Z}N(0,1)$ and $\bar{\eta}_{2}=a_{2}+b_{2}Z+c_{2}\sqrt{Z}N(0,1)$. Observe that $J_{1}(x)=F_{\bar{\eta}_{2}}^{(2)}(x)-F_{\bar{\eta}_{2}}^{(2)}(x)$, which can be seen from (13) when $k=2$. Due to the condition (30), we can assume either $a_{1}>a_{2},b_{1}<b_{2}$ or the otherway. Because if $a_{1}=a_{2},b_{1}=b_{2}$, then we clearly have $J_{1}(x)=0$. We can write $b_{2}=b_{1}+\tau$ with the understanding that if $\tau>0$ then $b_{2}>b_{1}$ and if $\tau<0$ then $b_{1}>b_{2}$. With this and from (30) we have $a_{2}=a_{1}-\tau EZ$. We conclude that $\bar{\eta}_{2}=(a_{1}-\tau EZ)+(b_{1}+\tau)Z+c_{2}\sqrt{Z}N(0,1)=\bar{\eta}_{1}-\tau(EZ-Z)$. From (13) when $k=2$, we have
$F_{\bar{\eta}_{2}}(x)=\int_{m_{1}}^{m_{2}}c_{2}\sqrt{z}\phi^{(2)}(\frac{x-a_{1}-b_{1}z}{c_{2}\sqrt{z}}+\tau\frac{EZ-z}{c_{2}\sqrt{z}})f(z)dz.$
Therefore we have
$\begin{split}J_{1}(x)=&\int_{m_{1}}^{m_{2}}c_{2}\sqrt{z}[\phi^{(2)}(\frac{x-a_{1}-b_{1}z}{c_{2}\sqrt{z}}+\tau\frac{EZ-z}{c_{2}\sqrt{z}})-\phi^{(2)}(\frac{x-a_{1}-b_{1}z}{c_{2}\sqrt{z}})]f(z)dz\\ =&\underbrace{\int_{m}^{m_{2}}c_{2}\sqrt{z}[\phi^{(2)}(\frac{x-a_{1}-b_{1}z}{c_{2}\sqrt{z}}+\tau\frac{EZ-z}{c_{2}\sqrt{z}})-\phi^{(2)}(\frac{x-a_{1}-b_{1}z}{c_{2}\sqrt{z}})]f(z)dz}_{H_{1}(x)}\\ +&\underbrace{\int_{m_{1}}^{m}c_{2}\sqrt{z}[\phi^{(2)}(\frac{x-a_{1}-b_{1}z}{c_{2}\sqrt{z}}+\tau\frac{EZ-z}{c_{2}\sqrt{z}})-\phi^{(2)}(\frac{x-a_{1}-b_{1}z}{c_{2}\sqrt{z}})]f(z)dz}_{H_{2}(x)}.\end{split}$ (31)
If $\tau>0$ (which is the case $b_{2}>b_{1}$), then $\tau\frac{EZ-z}{c_{2}\sqrt{z}}<0$ when $z\in[m,m_{2}]$ and $\tau\frac{EZ-z}{c_{2}\sqrt{z}}>0$ when $z\in[m_{1},m]$. Then by Lemma 3.9 we have
$\begin{split}H_{1}(x)\geq&\int_{m}^{m_{2}}c_{2}\sqrt{z}[\tau\frac{EZ-z}{c_{2}\sqrt{z}}]f(z)dz=\tau\int_{m}^{m_{2}}(EZ-z)f(z)dz\\ H_{2}(x)\geq&\tau\int_{m_{1}}^{m}(EZ-z)f(z)dz\end{split}$ (32)
Therefore when $\tau>0$ we have
$J_{1}(x)=H_{1}(x)+H_{2}(x)\geq\tau\int_{m_{1}}^{m_{2}}(EZ-z)f(z)dz=0.$
Now if $\tau<0$ (which is the case $b_{1}>b_{2}$), we have $\tau\frac{EZ-z}{c_{2}\sqrt{z}}>0$ when $z\in[m,m_{2}]$ and $\tau\frac{EZ-z}{c_{2}\sqrt{z}}<0$ when $z\in[m_{1},m]$. Again by using Lemma 3.9 above we obtain $J_{1}(x)\geq 0$.
From these two steps we conclude that $I_{2}(x)-I_{1}(x)=J_{2}(x)+J_{1}(x)\geq J_{2}(x)>0$ for all $x\in\mathbb{R}$ as $c_{1}<c_{2}$. This ends the proof. ∎
Lemma 3.11.
Let $\eta_{1}$ and $\eta_{2}$ be defined as in Lemma 3.10 above, with the same boundedness assumption on $Z$. Suppose that either
$a_{1}+b_{1}EZ>a_{2}+b_{2}EZ,\;c_{1}\leq c_{2},$
or
$a_{1}+b_{1}EZ\geq a_{2}+b_{2}EZ,\;c_{1}<c_{2}.$
Then $\eta_{1}\succ_{(2)}\eta_{2}$.
Proof.
First consider the case $a_{1}+b_{1}EZ>a_{2}+b_{2}EZ,\;c_{1}\leq c_{2}$. Set $\delta=(a_{1}+b_{1}EZ)-(a_{2}+b_{2}EZ)$ and define $\eta_{\delta}=\eta_{2}+\delta$. Then $\eta_{\delta}$ strictly first-order stochastically dominates $\eta_{2}$, i.e., $\eta_{\delta}\succ_{(1)}\eta_{2}$. Equivalently, $F_{\eta_{\delta}}^{(2)}(x)<F_{\eta_{2}}^{(2)}(x)$ for all $x$. Note also that $E\eta_{\delta}=E\eta_{1}$ and $c_{1}\leq c_{2}$. Define $I_{\delta}(x)$ and $I_{\eta_{1}}(x)$ analogously to the proof of Lemma 3.10. Repeating the same type of argument as there, we obtain $I_{\delta}(x)\geq I_{\eta_{1}}(x)$ for all $x$, which implies $F_{\eta_{\delta}}^{(2)}(x)\geq F_{\eta_{1}}^{(2)}(x)$ for all $x$. Combining these inequalities yields $F_{\eta_{2}}^{(2)}(x)>F_{\eta_{1}}^{(2)}(x)$ for every $x$.
Now consider the case $a_{1}+b_{1}EZ\geq a_{2}+b_{2}EZ,\;c_{1}<c_{2}$. Define $\delta$ and $\eta_{\delta}$ as above. Then clearly $F^{(2)}_{\eta_{\delta}}(x)\leq F^{(2)}_{\eta_{2}}(x)$ for all $x$ (note that here we may have $\delta=0$). Moreover, $E\eta_{\delta}=E\eta_{1}$ and $c_{1}<c_{2}$, so by Lemma 3.10 we have $F^{(2)}_{\eta_{\delta}}(x)>F^{(2)}_{\eta_{1}}(x)$ for all $x$. Hence, again we conclude $F_{\eta_{2}}^{(2)}(x)>F_{\eta_{1}}^{(2)}(x)$ for all $x$. ∎
The following Proposition extends the conclusions of the preceding Lemmas to encompass arbitrary mixing distributions $Z$ within the NMVM framework.
Proposition 3.12.
Fix any positive integer $k\geq 1$. Assume $Z\in L^{k}$. Consider the following two NMVM models $\eta_{1}=a_{1}+b_{1}Z+c_{1}\sqrt{Z}N$ and $\eta_{2}=a_{2}+b_{2}Z+c_{2}\sqrt{Z}N$, where $a_{1},a_{2},b_{1},b_{2}$ are any real numbers and $c_{1}>0,\;c_{2}>0$. Then if
$a_{1}+b_{1}EZ\geq a_{2}+b_{2}EZ\;\;\mbox{and}\;\;c_{1}>c_{2},$ (33)
or
$a_{1}+b_{1}EZ>a_{2}+b_{2}EZ\;\;\mbox{and}\;\;c_{1}\geq c_{2},$ (34)
we have $\eta_{1}\succ_{(2)}\eta_{2}$.
Proof.
We first prove the claim under the condition (33). Let $\delta=(a_{1}+b_{1}EZ)-(a_{2}+b_{2}EZ)$ (this number could be zero). Define $\eta_{\delta}=\delta+\eta_{2}$. It is clear that $F^{(2)}_{\eta_{\delta}}(x)\leq F^{(2)}_{\eta_{2}}(x)$ for all $x$. We have $E\eta_{\delta}=E\eta_{1}$ while $c_{1}<c_{2}$. If we can show that $\eta_{1}\succeq_{(2)}\eta_{\delta}$ then we are done. So everything boils down to showing $\eta_{1}\succeq_{(2)}\eta_{2}$ under the condition $a_{1}+b_{1}EZ=a_{2}+b_{2}EZ$ in (33). For each positive integer $m\geq 1$ define $Z_{m}=Z1_{\{1/m,m\}}(Z)+1/m1_{\{(0,1/m]\}}(Z)+m1_{\{m,\infty\}}(Z)$. Then clearly $Z^{m}$ is strictly positive and $Z_{m}\rightarrow Z$ in $L^{k}$. Also $EZ_{m}\rightarrow EZ$. Define $\eta_{1m}=a_{1}+b_{1}Z_{m}+c_{1}\sqrt{Z_{m}}N$ and $\eta_{2m}=a_{2}+b_{2}Z_{m}+c_{2}\sqrt{Z_{m}}N$. Let $\delta_{m}=(a_{1}-a_{2})+(b_{1}-b_{2})EZ_{m}$. Due to the condition $a_{1}+b_{1}EZ=a_{2}+b_{2}EZ$ we have $\delta_{m}\rightarrow 0$. Define $\bar{\eta}_{1m}=\delta_{m}+\eta_{2m}$. We have $F_{\bar{\eta}_{1m}}^{(2)}(x)\rightarrow F^{(2)}_{\eta_{2m}}(x)\rightarrow F^{(2)}_{\eta_{2}}(x)$ for each fixed point $x$ when $m\rightarrow\infty$ as $\delta_{m}\rightarrow 0$. Now define $\bar{I}_{1m}(x)=E(x-\bar{\eta}_{1m})^{+}$ and $I_{1m}(x)=E(x-\eta_{1m})^{+}$. We have $I_{1m}=E(x-\eta_{1m})^{+}\rightarrow E(x-\eta_{1})^{+}=F_{\eta_{1}}^{(2)}(x)$ as $m\rightarrow\infty$. Also we have $\bar{I}_{1m}(x)=F_{\eta_{1m}}^{(2)}(x-\delta_{m})\rightarrow F^{(2)}_{\eta_{2}}(x)$. Now, from the proof of Lemma 3.10, we have
$\bar{I}_{1m}(x)-I_{1m}(x)\geq(c_{2}-c_{1})\int_{1/m}^{m}z\phi^{(0)}(\frac{x-a_{1}-b_{1}z}{c_{1}\sqrt{z}})f_{m}(z)dz.$
The right-hand-side of this expression goes to $(c_{2}-c_{1})\int_{0}^{\infty}z\phi^{(0)}(\frac{x-a_{1}-b_{1}z}{c_{1}\sqrt{z}})f(z)dz>0$ while the left-hand-side goes to $F_{\eta_{2}}(x)-F_{\eta_{1}}(x)$.
Next we prove the claim under the condition (34). We let $\delta=(a_{1}-a_{2})+(b_{1}-b_{2})EZ>0$. Define $\eta_{\delta}=\delta+\eta_{2}$. Clearly $F^{(2)}_{\eta_{\delta}}(x)<F^{(2)}_{\eta_{2}}(x)$ for all $x$. Observe that $E\eta_{\delta}=E\eta_{1}$ while $c_{1}\geq c_{2}$. We need to show $F^{(2)}_{\eta_{\delta}}(x)\geq F^{(2)}_{\eta_{1}}(x)$ and this can be done by defining $Z_{m}$ as in the first step and by applying the method in the proof of Lemma 3.10. ∎
One of the main uses of these results appears in the following corollary.
Corollary 3.13.
Consider the model (1) and assume that $EZ<\infty$ and, in addition, that $\mu+\gamma EZ\neq 0$. Take the hyperplane $\omega^{T}(\mu+\gamma EZ)=c$ for an arbitrary real constant $c$, and let $\omega_{1}$ and $\omega_{2}$ be any two points lying on this hyperplane. If
$\omega_{1}^{T}\Sigma\omega_{1}<\omega_{2}^{T}\Sigma\omega_{2},$
then it follows that $(\omega_{1}^{T}X)\succ_{(2)}(\omega_{2}^{T}X)$.
Proof.
Note first that $\omega_{1}^{T}X\overset{d}{=}\omega_{1}^{T}\mu+\omega_{1}^{T}\gamma Z+\sqrt{\omega_{1}^{T}\Sigma\omega_{1}}N(0,1)$ and $\omega_{2}^{T}X\overset{d}{=}\omega_{2}^{T}\mu+\omega_{2}^{T}\gamma Z+\sqrt{\omega_{2}^{T}\Sigma\omega_{2}}N(0,1)$. The claim then follows directly from Proposition (3.12). ∎
As an application of the above results, we now present the following example.
Example 3.14.
Let $R$ denote the return vector of $d$ risky assets and suppose that
$R\sim GH_{d}(\lambda,\alpha,\beta,\delta,\mu,\Sigma),$
that is, $R$ follows a multivariate generalized hyperbolic distribution. For the definition of this distribution and the admissible parameter ranges, we refer to Chapter 2 of [15]. As stated in the first paragraph on page 78 of [15], $R$ admits the NMVM representation
$R\overset{d}{=}\mu+Z\Sigma\beta+\sqrt{Z}AN_{d},$
[…]
We note that the conclusion of Example 3.14 above continues to hold for any risk measure that is both law-invariant and SSD-consistent. As an illustration, consider the following family of risk measures, which satisfy these requirements:
$\hat{\rho}(X)=\sup_{\mu\in\mathcal{M}}\int_{(0,1]}CVaR_{\beta}(X)\mu(d\beta),$ (35)
where $\mathcal{M}$ denotes an arbitrary subset of $\mathcal{M}_{1}((0,1])$, the collection of probability measures on $(0,1]$. For additional details on this class of risk measures, we refer to Section 5.1 and the preceding sections of [12]. Furthermore, any finite-valued law-invariant coherent risk measure admits a representation of the above form; see, for instance, the paragraph preceding Corollary 5.1 in [12].
4 Applications
In this section, we examine the capital asset pricing model (CAPM) within our framework. We begin by establishing the following key lemma.
Lemma 4.1.
Consider two NMVM random variables $\eta_{1}=a_{1}+b_{1}Z+c\sqrt{Z}N(0,1)$ and $\eta_{2}=a_{2}+b_{2}Z+c\sqrt{Z}N(0,1)$ with $a_{1},a_{2},b_{1},b_{2}\in\mathbb{R}$ and $c>0$. Assume $EZ<\infty$ and $a_{1}+b_{1}EZ=a_{2}+b_{2}EZ$. Then
$F_{\eta_{1}}^{(2)}(x)=F_{\eta_{2}}^{(2)}(x),\;\;\forall x\in\mathbb{R}.$
Proof.
We define $Z_{m},\eta_{m1},\eta_{m2}$ as in the proof in Proposition (3.12). The corresponding second order CDFs are denoted by $F^{(2)}_{\eta_{1m}}$ and $F^{(2)}_{\eta_{2m}}$. From the proof of Lemma 3.10 we have $F^{(2)}_{\eta_{2m}}(x)-F^{(1)}_{\eta_{2m}}(x)=J_{m1}(x)$, where $J_{m1}(x)$ is the $J_{1}(x)$ that corresponds to $\eta_{m1},\eta_{m2}$. By following the same argument in Step 2 of this Lemma we can show that $F^{(2)}_{\eta_{2m}}-F^{(1)}_{\eta_{2m}}=J_{m1}(x)\geq 0$. Now we switch the positions of $F^{(2)}_{\eta_{1m}}(x)$ and $F^{(2)}_{\eta_{2m}}(x)$ and obtain $F^{(1)}_{\eta_{2m}}-F^{(1)}_{\eta_{2m}}\geq 0$. These show that $F^{(1)}_{\eta_{2m}}(x)=F^{(1)}_{\eta_{2m}}(x)$. For the general case we take limit as $m\rightarrow\infty$. ∎
This Lemma establishes that $\eta_{1}\succeq_{(2)}\eta_{2}$ and, at the same time, $\eta_{2}\succeq_{(2)}\eta_{1}$ whenever $a_{1}+b_{1}EZ=a_{2}+b_{2}EZ$, independent of the specific values of $a_{1},a_{2},b_{1},$ and $b_{2}$. This observation has an important implication. Before presenting it, we briefly recall several relevant results from [27]. Let $F_{\eta}^{(-1)}(p)=\inf\{x:F_{\eta}(x)\geq p\},\;0<p\leq 1$ denote the left-continuous inverse of the CDF of $\eta$. Define
$F_{\eta}^{(-2)}(p)=\int_{0}^{p}F^{(-1)}_{\eta}(\alpha)\,d\alpha,\;\;0<p\leq 1.$
Theorem 3.1 in [27] shows that $F_{\eta}^{(-2)}=[F^{(2)}_{\eta}$$^{\star}$, i.e., the conjugate of $F^{(2)}_{\eta}$. Consequently, we immediately have $F_{\eta_{1}}^{(-2)}=F_{\eta_{2}}^{(-2)}$ whenever $F^{(2)}_{\eta_{1}}=F^{(2)}_{\eta_{2}}$. The same paper also establishes that $CVaR_{\alpha}=\frac{F_{\eta}^{(-2)}(p)}{p}$ with $\alpha=1-p$. We will make use of this relation in the proof of the following Proposition.
Proposition 4.2.
Let $\eta=a+bZ+c\sqrt{Z}N$ be an arbitrary real-valued NMVM model, and let $\rho$ be any law-invariant coherent risk measure satisfying $\rho(0)=0$. Then
$\rho(\eta)=a+bEZ+c\rho(\sqrt{Z}N).$
Consequently, for any portfolio $x$ we obtain
$\rho(\omega^{T}X)=\omega^{T}\mu+\omega^{T}\gamma EZ+\sqrt{\omega^{T}\Sigma\omega}\rho(\sqrt{Z}N).$ (36)
Proof.
By Lemma 4.1 and the discussion preceding this proposition, we know that $CVaR_{\alpha}(\eta)=CVaR_{\alpha}(\bar{\eta})$ for any $\bar{\eta}=\bar{a}+\bar{b}Z+\bar{c}\sqrt{Z}N$ such that $a+bEZ=\bar{a}+\bar{b}EZ$. Note that we can write
$CVaR_{\alpha}(\eta)=E\eta+CVaR_{\alpha}(\eta-E\eta).$
The random variable $\eta-E\eta$ is again an NMVM model, now with zero mean, and thus its risk coincides with that of $c\sqrt{Z}N$. Hence
$CVaR_{\alpha}(\eta)=a+bEZ+CVaR_{\alpha}(c\sqrt{Z}N).$
By the positive homogeneity of $CVaR$, this yields the stated formula in the case of $CVaR$. For a general law-invariant convex risk measure $\rho$, the claim follows by applying relation (35). ∎
Remark 4.3.
The relation (36) above implies that, for any law-invariant convex risk measure $\rho$ (such as Expected Shortfall / CVaR), the risk of a mixture asset can be decomposed into two parts: the expectation of the linear term $\omega^{T}\mu+\omega^{T}\gamma EZ$ and the risk of the purely normal component, multiplied by the factor $\sqrt{\omega^{T}\Sigma\omega}$. This representation turns an otherwise intricate, distribution-dependent portfolio risk evaluation into a more tractable expression.
Tangent portfolio: The tangent portfolio is designed to achieve the best possible balance between risk and return. When a risk-free asset is available, it becomes the single optimal risky portfolio for every investor, no matter how risk-averse they are. Among all feasible portfolios composed solely of risky assets, the tangent portfolio attains the maximum Sharpe Ratio. Within the framework of the Capital Asset Pricing Model (CAPM), the tangent portfolio acquires an even more prominent interpretation: it corresponds to the Market Portfolio.
To derive the tangent portfolio in our setting, we consider a scenario where investors can allocate their wealth across $d$ risky assets, whose return vector is specified in (1), and a risk-free asset with return $r_{f}$. For a portfolio $\omega\in\mathbb{R}^{d}$, the resulting portfolio return is
$R_{\omega}=\omega^{T}X+(1-\omega^{T}\mathbf{1})r_{f}.$
By Proposition 4.2, the associated risk is
$\begin{split}\rho(R_{\omega})=&\omega^{T}(\mu+\gamma EZ)+\sqrt{\omega^{T}\Sigma\omega}\rho(\sqrt{Z}N)+(1-\omega^{T}\mathbf{1})r_{f}\\ =&r_{f}+\omega^{T}(\mu-\mathbf{1}r_{f}+\gamma EZ)+\sqrt{\omega^{T}\Sigma\omega}\rho(\sqrt{Z}N).\end{split}$ (37)
For brevity, we henceforth set $\rho_{0}:=\rho(\sqrt{Z}N)$. In the next result, we identify the corresponding tangent portfolio.
Proposition 4.4.
The market portfolio in this setting is
$\omega_{m}=\frac{1}{\mathbf{1}^{T}\Sigma^{-1}(\mu-\mathbf{1}r_{f}+\gamma EZ)}\Sigma^{-1}(\mu-\mathbf{1}r_{f}+\gamma EZ).$
The associated market return can be written as
$R_{m}=\omega_{m}^{T}X=u_{m}+\beta_{m}Z+\sigma_{m}\sqrt{Z}N,$
where
$u_{m}=\omega_{m}^{T}\mu,\;\beta_{m}=\omega_{m}^{T}\gamma,\;\sigma_{m}=\sqrt{\omega_{m}^{T}\Sigma\omega_{m}}.$
Moreover,
$\rho(R_{m})=u_{m}+\gamma_{m}EZ+\sigma_{m}\rho_{0}.$
Proof.
We maximize the Sharpe ratio
$\frac{ER_{\omega}-r_{f}}{\rho(R_{\omega})}=\frac{\omega^{T}(\mu-\mathbf{1}r_{f}+\gamma EZ)}{\rho(R_{\omega})}.$
First note that, from (37) we have
$\frac{d\rho(R_{\omega})}{d\omega}=(\mu-\mathbf{1}r_{f}+\gamma EZ)+\frac{\rho_{0}}{2\sqrt{\omega^{T}\Sigma\omega}}\Sigma\omega.$
Imposing the first-order condition yields that the optimal portfolio $\omega^{\star}$ is of the form
$\omega^{\star}=L(\omega^{\star})\Sigma^{-1}(\mu-\mathbf{1}r_{f}+\gamma EZ),$
for some function $L(\omega^{\star})$. Normalizing by the total weight, we obtain
$\omega_{m}=\frac{\omega^{\star}}{\mathbf{1}^{T}\omega^{\star}},$
which coincides with the expression stated in the proposition. The remaining assertions follow directly from the definitions. ∎
[…]
In what follows, we derive the corresponding capital asset pricing model (CAPM) based on the analysis developed in this paper. To do so, we begin by introducing the following risk measure:
$\bar{\sigma}(\eta)=\frac{\rho(\eta)-E\eta}{\rho(\sqrt{Z}N)}=\frac{\rho(\eta)-E\eta}{\rho_{0}},$ (38)
for any NMVM random variable $\eta=a+b\eta+c\sqrt{Z}N$. Due to Proposition 4.2 we have $\bar{\sigma}(\eta)=c$.
Theorem 4.5.
For any return $R=a+bZ+c\sqrt{Z}N$, we have
$ER-r_{f}=\frac{\rho(R)-ER}{\rho(R_{m})-ER_{m}}(ER_{m}-r_{f}).$ (39)
Proof.
We follow the idea of proof of the traditional CAPM. Consider a risky asset $i$ with return $R_{i}=a_{i}+b_{i}Z+c_{i}\sqrt{Z}N$ and the market return $R_{m}$. Invest a fraction $\omega$ of an investor’s wealth in asset $i$ and the remaining fraction $(1-\omega)$ in the market portfolio. The corresponding return is denoted by $R_{\omega}=\omega R_{i}+(1-\omega)R_{m}$. We have $ER_{\omega}=\omega ER_{i}+(1-\omega)ER_{M}$ and $\bar{\sigma}(R_{\omega})=\omega c_{i}+(1-\omega)\sigma_{m}$. We calculate $\frac{dER_{\omega}}{d\omega}=ER_{i}-ER_{m}$ and $\frac{d\bar{\sigma}}{d\omega}=c_{i}-\sigma_{m}$. We have $\frac{dER_{\omega}}{d\omega}|_{\omega=0}=ER_{i}-ER_{m}$ and $\frac{d\bar{\sigma}}{d\omega}|_{\omega=0}=c_{i}-\sigma_{m}$. We utilize the relation
$\frac{dER_{\omega}}{d\omega}|_{\omega=0}]/[\frac{d\bar{\sigma}}{d\omega}|_{\omega=0}]=\frac{ER_{m}-r_{f}}{\sigma_{m}}.$
This leads into the equation
$\frac{ER_{i}-ER_{m}}{c_{i}-\sigma_{m}}=\frac{ER_{m}-r_{f}}{\sigma_{m}}.$
After some algebra we obtain
$ER-r_{f}=\frac{c_{i}}{\sigma_{m}}(ER_{m}-r_{f}).$
Observe that, due to Proposition 4.2, we have $c_{i}=\frac{\rho(R_{i})-ER_{i}}{\rho_{0}}$ and $\sigma_{m}=\frac{\rho(R_{m})-ER_{m}}{\rho_{0}}$. This ends the proof. ∎
Observe that we always have $\rho(R)-ER>0$ and $\rho(R_{m})-ER_{m}>0$, because $\sqrt{Z}N$ is a symmetric random variable, which guarantees that $\rho(\sqrt{Z}N)>0$ in (36). Since, as shown in (40), the tangent portfolio $\omega_{m}$ is derived using both the mean and the risk measure (38), the computations in the proof of Theorem 4.5 proceed analogously to the derivation of the classical CAPM formula. Furthermore, the tangent portfolio in Proposition 4.4 can likewise be characterized as the solution to the maximization of the Sharpe ratio
$\frac{ER_{\omega}-r_{f}}{\bar{\sigma}(R_{\omega})},$
where the risk measure in (38) is employed.
Remark 4.6.
Recall that the standard CAPM is
$ER-r_{f}=\frac{Cov(R,R_{m})}{Var{R_{m}}}(ER_{m}-r_{f}),$
in which the covariance between the asset return $R$ and the market return $R_{m}$ enters explicitly. By contrast, the skewness-induced CAPM (SI-CAPM) (39) does not feature this covariance term. Nonetheless, under our specification $R=a+bZ+c\sqrt{Z}N$, the return $R$ remains correlated with $R_{m}$, and in fact
$Cov(R_{m},R)=b\beta_{m}Var(Z)+c\sigma_{m}EZ.$
The classical CAPM relies on normally distributed returns and takes variance as the relevant notion of risk (which leads to beta as the risk measure). Our relation (39) yields an analogous linear link between expected return and risk. Even when returns follow a heavy-tailed mixture distribution (such as a t-distribution) and risk is assessed via a more intricate measure (such as Expected Shortfall), the resulting equilibrium pricing relation preserves the simple linear form
$ER=r_{f}+\beta(ER_{m}-r_{f}).$
However, in this framework, the beta $\beta$ is determined by the parameters of the mixture model rather than solely by the covariance with the market.
Remark 4.7.
The risk measure $\bar{\sigma}(\cdot)$ introduced in (38) is not convex. It can instead be interpreted as a risk measure analogous to the usual standard deviation. Under this risk measure, we clearly obtain $\bar{\sigma}(R_{\omega})=\sqrt{\omega^{T}\Sigma\omega}$ and $\bar{\sigma}^{2}(R_{\omega})=\omega^{T}\Sigma\omega$ (see also (37)). The tangent portfolio in Proposition 4.4 can likewise be derived using the risk measure $\bar{\sigma}^{2}(\cdot)$. The derivation proceeds in the same way as the construction of the tangent portfolio in the Markowitz setting, thereby providing additional justification for the computations carried out in Theorem 4.5.
Remark 4.8.
The formula (39) can be equivalently expressed as
$\frac{ER-r_{f}}{\rho(R)-ER}=\frac{ER_{m}-r_{f}}{\rho(R_{m})-ER_{m}},$
which shows that the ratio of the expected excess return $ER-r_{f}$ of any asset to its excess risk (over the expected return) $\rho(R)-ER$ is equal to the corresponding ratio for the market portfolio.
Mean-$\rho$ frontier curve: Next, we write down the corresponding frontier curve. For this, we consider investments only on risky assets (1). The return $R_{\omega}$ of a portfolio $\omega$ with $\omega^{T}\mathbf{1}=1$ is given by $R_{\omega}=\omega^{T}\mu+\omega^{T}\gamma Z+\sqrt{\omega^{T}\Sigma\omega}\sqrt{Z}N$. Since $\rho(R_{\omega})=\omega^{T}(\mu+\gamma EZ)+\sqrt{\omega^{T}\Sigma\omega}\rho_{0}$ by Proposition 4.2, the frontier portfolios are obtained by solving the following optimization problem
[…]
for all return levels $r$. We introduce the following notations:
$\mathcal{A}=(\mu+\gamma EZ)^{T}\Sigma^{-1}(\mu+\gamma EZ),\;\mathcal{B}=(\mu+\gamma EZ)^{T}\Sigma^{-1}\mathbf{1},\;\mathcal{C}=\mathbf{1}^{T}\Sigma^{-1}\mathbf{1},\;\triangle=\mathcal{A}\mathcal{C}-\mathcal{B}.$
The solution of problem (40) is given by
$\omega_{r}=g+rh,$
where $g=\frac{1}{\triangle}[\mathcal{A}(\Sigma^{-1}\mathbf{1})-\mathcal{B}(\Sigma^{-1}\mu)]$ and $h=\frac{1}{\triangle}[\mathcal{C}(\Sigma^{-1}\mu)-\mathcal{B}(\Sigma^{-1}\mathbf{1})]$. One can easily calculate that
$\omega^{T}_{r}\Sigma\omega_{r}=\frac{1}{\triangle}(\mathcal{C}r^{2}-2\mathcal{B}r+\mathcal{A}).$
We know that
$\rho(R_{\omega_{r}})=\omega_{r}^{T}(\mu+\gamma EZ)+\sqrt{\omega_{r}^{T}\Sigma\omega_{r}}\,\rho_{0},$
and, by construction, $\omega_{r}^{T}(\mu+\gamma EZ)=r$. Therefore, the frontier can be written in closed form as
$\rho=r+\frac{\rho_{0}}{\sqrt{\triangle}}\sqrt{\mathcal{C}r^{2}-2\mathcal{B}r+\mathcal{A}},$ (41)
[…]
We emphasize here that the mean-$\rho$ frontier (41) is a powerful and adaptable tool, generalizing the traditional mean–variance framework to work with any risk measure $\rho$. In particular, it can reflect investor-specific notions of risk, deal with non-normal, heavy-tailed, and skewed return distributions, emphasize tail risk for regulatory or conservative investment objectives, and offer a solid theoretical basis for diversification benefits even when variance is not the chosen risk metric. By replacing a single, fixed risk measure with a generic $\rho$, the frontier evolves into a flexible framework suited to contemporary portfolio management, where risk is multi-dimensional and return distributions are typically non-normal.
Globally minimum risk portfolio: The global minimum risk (GMR) portfolio is of substantial practical and theoretical relevance in finance. Its significance arises from its distinctive characteristics and its function in both portfolio design and financial economics. By construction, it is the risk-minimizing portfolio, and therefore the one whose future value an investor can forecast with the greatest degree of certainty. Whereas every other portfolio on the efficient frontier entails a balance between risk and expected return, the GMR portfolio is the one that most effectively reduces uncertainty. For an investor with strong risk aversion, this portfolio is therefore optimal. The global minimum risk portfolio is obtained by solving the following optimization problem:
[…]
We denote the optimizer of (42) by $\omega_{GMR}$. By Proposition 4.2, we have $\rho(\omega^{T}X)=\omega^{T}(\mu+\gamma EZ)+\sqrt{\omega^{T}\Sigma\omega}\rho_{0}$. Consequently, problem (42) reduces to an optimization problem involving functionals of quadratic forms, a type of problem analyzed in [19]. We decompose
$\Sigma=\begin{pmatrix}\Sigma_{11}&\Sigma_{12}\\ \Sigma_{21}&\sigma_{dd}\end{pmatrix}$,
where $\Sigma_{11}$ is a $(d-1)\times(d-1)$ matrix and $\sigma_{dd}$ is the $(d,d)$-entry of $\Sigma$. Let $\mathbf{1}_{1}$ denote the (d1)(d-1)-dimensional column vector of ones. Define $m_{1}=(\mu_{1}+\gamma_{1}EZ,\cdots,\mu_{d-1}+\gamma_{d-1}EZ)^{T}$ and $m_{2}=\mu_{d}+\gamma_{d}EZ$ (so that $\mu+\gamma EZ=(m_{1},m_{2})$). Set
$D=m_{2}\mathbf{1}_{1}-m_{1},\;\;Q=\Sigma_{11}-\mathbf{1}_{1}\Sigma_{12}^{T}-\Sigma_{12}\mathbf{1}^{T}+\sigma_{dd}\mathbf{1}_{1}\mathbf{1}_{1}^{T}.$
Then, provided that $\rho_{0}>\sqrt{D^{T}Q^{-1}D}$ (see Theorem 1 in [19]), we obtain
$\omega_{GMR}=\frac{1}{\mathbf{1}^{T}\Sigma^{-1}\mathbf{1}}\Sigma^{-1}\mathbf{1}-\frac{1}{\sqrt{(\rho_{0}^{2}-D^{T}Q^{-1}D)(\mathbf{1}^{T}\Sigma\mathbf{1})}}(D^{T}Q^{-1},\mathbf{1}_{1}^{T}Q^{-1}D)^{T}.$
Acknowledgments
The author wishes to express gratitude to Alexander Schied for his comments and guidance during the initial phases of this project.